ANALYSIS OF THE A^{2}Π STATES OF Sr^{79}Br AND Sr^{81}Br: DEPERTURBATION OF THE FIRST EXCITED STATE

\underline{C. S. DICKINSON} and J. A. COXON, Department of Chemistry, Dalhousie University, Halifax, N.S.

Canada B3H 4J3.

High-resolution laser excitation spectra of the A^{2}Π- X^{2}Σ^{+} system have been obtained for the 0-0 through 3-0 bands of the two principal isotopomers of SrBr. The molecules were produced by reacting stontium vapor with methyl bromide in a Broida oven. A Coherent 699- 29 ring dye laser was used as a probe for both selective detection and resolved fluorescence. Perturbations of the A^{2}Π_{1/2} state with the closely lying B^{2}Σ^{+} state, (∼650cm^{−1} higher in energy) are

being characterized. An analysis of the ”crossing” of the v=3 level of the A state with the v=0 level of the B state is required in order to obtain physically significant parameters.